A rectangle has a perimeter of 96cm. If the length is 2 times it’s width, what are the dimensions of the rectangle?
Answer by ikleyn(53266)

.
The perimeter = x + 2x + x + 2x = 96.

6x = 96

x = 96/6 = 16.

Answer. The dimensions of the rectangle are 16 cm (width) and 32 cm (length).